# Resource Allocation via Backscatter-Aware Transmit Antenna Selection for Low-PAPR and Ultra-Reliable WSNs

### v1

@@ V-A Simulation Setup
Parameter Category Parameter Value
System and Channel Model
System Scheme Proposed BC-TAS MD-OFDM
Baseline Schemes Conventional TAS / MMSE $N_{T}=4,N_{R}=4$
Standard Context IEEE 802.11be EHT Physical Layer
Channel Model Indoor Multipath Fading TGn Models A–F
Fading Distribution Rayleigh Independent Fading Paths
CSI Assumption P-CSI / I-CSI Perfect / Estimation Error $\sigma_{e}^{2}$
OFDM and PHY Layer
Subcarrier Count ($N$) $N_{\mathrm{SC}}$ 256 / 1024
Cyclic Prefix Length ($N_{\mathrm{CP}}$) $L_{\mathrm{CP}}$ 32 samples
Modulation Scheme $M$-QAM 16-QAM / 64-QAM
Total Simulated Bits $N_{\mathrm{bits}}$ $10^{7}$ bits
Non-Linearity and Field Shaping
PA Model Rapp Model $p=2$ (Smoothness factor)
MOFS Weights $\lambda_{T},\lambda_{V}$ $[0.1,0.5,1.0]$
Tag/Victim Distances $d_{L},d_{T},d_{V}$ 10 m, 2 m, 5 m
Kalman Process Noise ($Q$) $Q_{KF}$ $1\times 10^{-4}$
Kalman Meas. Noise ($R$) $R_{KF}$ $1\times 10^{-2}$
Monte Carlo Trials $N_{trials}$ 1,000
V-B Comparative Analysis with State-of-the-Art (SOTA) Baselines
To validate the competitive standing of the proposed BC-TAS framework, we benchmarked its performance against four distinct SOTA categories: fixed-diversity baselines [SOTA_2_MU_STBC], low-complexity selection methods [SOTA_1_MMSE], theoretical upper bounds (OES) [SOTA_3_Exhaustive_Search], and backscatter-specific multi-antenna contexts [SOTA_4_Backscatter_Multi_Antenna].
@@ V-B Comparative Analysis with State-of-the-Art (SOTA) Baselines
As shown in the simulation results in Fig. 1, the proposed BC-TAS significantly outperforms the industry-standard Norm-Based Antenna Selection (NBAS) [SOTA_5_NBAS]. At an average SNR of 10 dB, BC-TAS achieves a BER of $\approx 3.2\times 10^{-4}$, whereas NBAS remains at $\approx 2.6\times 10^{-2}$. This order-of-magnitude improvement is attributed to our model’s ability to exploit frequency-selective diversity rather than relying on a static norm-based metric. The proposed scheme achieves a near-optimal slope while maintaining robustness to high CSI error ($20\%$).
V-C Impact of CSI Imperfection and Robustness Analysis
A critical requirement for practical backscatter deployments is resilience to CSI estimation errors, which are prevalent due to the low-power nature of the tag. We evaluate the robustness of the proposed BC-TAS framework under three regimes: Perfect CSI (P-CSI), 5% estimation error (csi_error_low=0.05csi\_error\_low=0.05), and a severe 20% error condition (csi_error_high=0.2csi\_error\_high=0.2).
As illustrated in Fig. 1, the BC-TAS (Perfect) configuration defines the heuristic’s lower bound, tracking closely with the MD-OFDM (Max Gain) ceiling until approximately 10 dB SNR. In this regime, the system effectively exploits the full spatial-frequency diversity available in the TGn Model F environment.
  • Moderate Imperfection (5% Error): Under a 5% CSI estimation error, the BC-TAS model exhibits a negligible SNR penalty of approximately 1.5 dB at a target BER of $10^{-3}$. Notably, the diversity slope remains steep, indicating that the multi-objective selection metric is still capable of identifying favorable subcarrier-antenna mappings despite minor phase and magnitude perturbations.
  • Severe Imperfection (20% Error): At 20% error, the system encounters a performance floor at a BER of $\approx 1.8\times 10^{-2}$. While the high error rate significantly degrades the selection accuracy, the proposed BC-TAS (20% Error) remarkably maintains a lower BER than the Norm-Based TAS and MMSE 4x4 baselines in the low-to-mid SNR regions (0–8 dB).
The distinct robustness of the proposed model over NBAS is attributed to the fundamental difference in their selection logic. While the performance of NBAS [SOTA_5_NBAS] collapses under severe 20% CSI error—often erroneously selecting a ”weak” antenna path due to noisy norm estimations—the multi-objective BC-TAS remains more resilient. By considering the average channel state across all 256 subcarriers through its heuristic metric, BC-TAS provides a ”smoothing” effect against noise. This prevents the selection of consistently poor paths, ensuring that the probability of selecting a deep-fade antenna across the entire bandwidth is significantly minimized, thereby maintaining a reliable communication link even under high CSI aging or estimation error.
V-D Computational Efficiency vs. Performance Ceiling
A critical finding of this study is the relationship between the proposed heuristic and the Optimal Exhaustive Search (OES) baseline. While OES [SOTA_3_Exhaustive_Search] defines the absolute performance ceiling—achieving zero BER at 8 dB—it requires a per-carrier exhaustive search across all $N_{sc}$ subcarriers.
[…]
@@ V-C Impact of CSI Imperfection and Robustness Analysis
V-E BER Performance and Environmental Robustness
The primary performance metric is the BER versus the SNR. The diversity gain provided by the BC-TAS algorithm is demonstrated by the steeper slope of the BER curve compared to the MMSE baseline.
@@ V-E BER Performance and Environmental Robustness
To validate the model’s robustness, performance is analyzed across the full suite of TGn channel models (A–F), as illustrated in Fig. 2. We observe that while the system maintains high reliability in all scenarios, the environment significantly impacts the diversity order. Model A (Flat/Small Office) exhibits the best performance due to the lack of severe frequency selectivity.
In contrast, as we move toward Model F (Large Open Space) with higher delay spreads, the increased frequency-selective fading introduces more frequent deep fades. However, the MD-OFDM selection diversity effectively mitigates these effects, achieving a $10^{-3}$ BER at approximately 12–14 dB SNR across all dispersive models. This demonstrates the framework’s ability to maintain a consistent sensing-communication link regardless of the physical indoor layout.
V-F Outage Probability and Diversity Gain Analysis
A critical bottleneck for WSNs is the prevalence of deep fading in indoor environments, which leads to link outages and energy-inefficient retransmissions. To evaluate the robustness of BC-TAS, we analyze the Cumulative Distribution Function (CDF) of the instantaneous SINR, $\gamma$, in a frequency-selective TGn Model D environment.
The outage probability is defined as $P_{\mathrm{out}}=P(\gamma<\gamma_{\mathrm{th}})$, where the target threshold is set to $\gamma_{\mathrm{th}}=7$ dB. This threshold represents the minimum operating SINR required for reliable QPSK demodulation and successful packet decoding in standard low-power protocols.
Figure 3: Outage Probability CDF in TGn Model D at an average $\mathrm{SNR}=16$ dB.
The simulation results, illustrated in Fig. 3 on a semi-logarithmic scale, reveal that the proposed BC-TAS scheme significantly reshapes the SINR distribution. While conventional random selection closely tracks the Theoretical Rayleigh (SISO) baseline—exhibiting an outage probability of approximately 0.10580.1058—the BC-TAS scheme suppresses the lower tail of the distribution through high-order selection diversity. The proposed BC-TAS scheme demonstrates a significantly steeper diversity slope compared to the Theoretical Rayleigh (SISO) and conventional random selection baselines. Specifically, at the $7$ dB threshold, BC-TAS achieves an outage probability of $0.000175$, representing a $604.77\times$ reduction in link outages.
@@ V-F Outage Probability and Diversity Gain Analysis
As detailed in Table III, this massive improvement stems from the transition from a diversity order of $L=1$ (SISO) to an achieved diversity order of $L=N_{T}=4$. The table demonstrates that the reliability gain grows exponentially as additional spatial paths are exploited. In Rayleigh fading, the outage probability for selection diversity scales as $(P_{\mathrm{out,SISO}})^{N_{T}}$; thus, the observed three-order-of-magnitude reduction is consistent with theoretical expectations for a four-antenna selection system. By effectively ”ignoring” subcarriers in deep fades, BC-TAS ensures that the signal remains above the decoding threshold almost $99.99\%$ of the time, providing a robust backscatter-aware link without increasing hardware complexity.
V-G Backscatter Crest Factor (BCF) and Link Stability
While BER and outage probability characterize the primary link, the stability of the reflected field at the passive tag is governed by the BCF. As derived in Lemma 1, the BC-TAS framework utilizes a field-flattening cost term to minimize the temporal peaks of the backscatter signal.
To ensure statistical rigor, we performed $1,000$ Monte Carlo realizations for each antenna configuration ($N_{t}\in\{2,4,8,16,32\}$). Furthermore, to mitigate the impact of imperfect CSI and prevent erratic antenna switching, we integrated a scalar Kalman Filter to smooth the estimated channel gains before selection.
@@ V-G Backscatter Crest Factor (BCF) and Link Stability
As illustrated in Fig. 4, the proposed BC-TAS (Kalman) model exhibits a superior downward trend in BCF compared to all baselines. At $N_{t}=32$, BC-TAS achieves a BCF of $\approx 14.0$ dB, representing a $1.1$ dB improvement over the OES and a $2.5$ dB improvement over the industry-standard NBAS. The proposed BC-TAS (Kalman) achieves a monotonic reduction in signal peaks, outperforming OES and SOTA baselines.
This gap is significant because OES [SOTA_3_Exhaustive_Search], while optimal for the primary receiver’s BER, is ”tag-blind” and selects antennas that create constructive interference peaks at the tag’s location. In contrast, BC-TAS intelligently sacrifices marginal SNR gains to maintain a ”flatter” reflected field. The inclusion of the Kalman filter is vital here; by providing a smoothed estimate of the channel state, it prevents the high BCF values associated with rapid, noise-driven antenna switching across adjacent subcarriers.
V-H PA Distortion Analysis
To evaluate the impact of non-linear hardware impairments, we employ the Rapp PA model with a smoothness factor $p=2$ and a saturation voltage $V_{sat}=1$. The distortion is quantified through Error Vector Magnitude (EVM) and spectral compliance with the IEEE 802.11n mask.
@@ V-H PA Distortion Analysis
As illustrated in the smoothed analysis in Fig. 5, the proposed BC-TAS maintains a consistent linearity advantage over the MMSE MU-MIMO baseline across the entire operational range. While the MMSE baseline remains significantly above the 16-QAM EVM limit of $5.6\%$ throughout the saturation-to-linear transition, the proposed BC-TAS exhibits a steep descent toward compliance as the IBO increases. Specifically, BC-TAS approaches the required $5.6\%$ threshold at an IBO of approximately $13$ dB, demonstrating a significant reduction in constellation distortion compared to MU-MIMO.
Furthermore, the indoor spectral compliance analysis in Fig. 8 confirms that the BC-TAS architecture successfully suppresses out-of-band emissions. Even at an IBO of $10$ dB, the normalized PSD remains strictly within the 802.11n spectral mask limits, with the selection diversity effectively mitigating the spectral regrowth typically associated with high-PAPR OFDM signals. By operating the PA with less back-off while maintaining signal integrity, BC-TAS provides a double benefit: it maximizes the average transmit power available for the backscatter tag’s energy harvesting while simultaneously ensuring robust coexistence with legacy primary receivers in the 2.4/5 GHz bands.
V-I Hardware Performance and Complexity Analysis
The hardware feasibility of the proposed BC-TAS is evaluated by analyzing its robustness against PA non-linearities and its overall energy footprint compared to conventional MU-MIMO baselines.
V-I1 PAPR Reduction and Statistical Analysis
The PAPR is a critical bottleneck for PA efficiency in multi-antenna systems. We evaluate the statistical distribution of these peak power events using the Complementary Cumulative Distribution Function (CCDF), which defines the probability that the PAPR of an OFDM symbol exceeds a specific threshold $\gamma_{0}$.
@@ V-I1 PAPR Reduction and Statistical Analysis
As illustrated in Fig. 6, the proposed BC-TAS architecture exhibits a profound statistical advantage over conventional multiplexing-based methods. While both systems converge at low PAPR thresholds where average signal behavior dominates, they diverge significantly in the high-power tail region. At a clipping probability of $10^{-2}$ (1%), the proposed BC-TAS maintains a PAPR of approximately 7.1 dB, whereas the conventional MMSE MU-MIMO baseline reaches 9.5 dB.
This significant 2.4 dB reduction in peak power is a direct consequence of the selection-based architecture. By leveraging the backscatter link to intelligently select subcarrier-antenna mappings that minimize constructive interference, BC-TAS effectively ”truncates” the probability of extreme peak events. This reduction in the PAPR tail is the primary driver for the improved hardware linearity and energy harvesting efficiency observed in subsequent analyses.
V-I2 Energy Efficiency (EE)
Energy Efficiency is defined as the ratio of spectral efficiency to the total power consumption, including RF chain activation and computational costs ($P_{RF}+P_{SEL}$).
@@ V-I2 Energy Efficiency (EE)
Our results in Fig. 7 demonstrate that BC-TAS provides a massive energy efficiency advantage, peaking at 0.075 bits/Hz/mW. In contrast, the MMSE baseline remains below 0.01 bits/Hz/mW due to the simultaneous activation of $N_{t}$ RF chains and high-complexity matrix inversions.
V-I3 Spectral Mask Compliance
Using the Rapp non-linearity model ($\rho=2$), we evaluate spectral regrowth against the IEEE 802.11n 20 MHz spectral mask.
Figure 8: Indoor Spectral Compliance: BC-TAS Transmit Spectrum under IEEE 802.11n Mask.
The normalized Power Spectral Density (PSD) in Fig. 8 confirms that BC-TAS is fully compliant with the 802.11n mask even at low IBO levels. Sideband emissions are suppressed below 45-45 dB, allowing the transmitter to operate closer to the PA’s saturation region to maximize harvestable power for backscatter tags.
V-J Joint Optimization and Fundamental Trade-off
The flexibility of the BC-TAS framework is showcased through the fundamental trade-off between communication reliability (Primary Link BER) and the interference suppression benefit ($\Delta I$) afforded to the victim receiver.

### v2

@@ VI-A Simulation Setup
Parameter Value / Configuration
System Scheme BC-TAS MD-OFDM ($N_{T},N_{R}=4$)
Standard Context IEEE 802.11 TGn (EHT PHY)
Channel Model TGn Models A–F (Rayleigh)
CSI Assumption Perfect / $\sigma_{e}^{2}$ Error (5%, 20%)
Subcarriers ($N_{\mathrm{SC}}$) 256 (80 MHz BW)
Cyclic Prefix ($L_{\mathrm{CP}}$) 32 samples
Modulation ($M$-QAM) 16-QAM / 64-QAM
Simulated Bits ($N_{\mathrm{bits}}$) $10^{7}$ bits
PA Model (Rapp) $p=2$, $V_{sat}=1$
MOFS Weights $\lambda_{T},\lambda_{V}\in\{0.1,0.5,1.0\}$
Distances ($d_{L},d_{T},d_{V}$) 10 m, 2 m, 5 m
Kalman ($Q_{KF},R_{KF}$) $10^{-4},10^{-2}$
Monte Carlo Trials 1,000
VI-B Comparative Analysis with SOTA Baselines
To validate the BC-TAS framework, we benchmark it against four SOTA categories: fixed-diversity baselines [SOTA_2_MU_STBC], low-complexity selection [SOTA_1_MMSE], theoretical upper bounds optimal exhaustive search (OES) [SOTA_3_Exhaustive_Search], and backscatter-specific multi-antenna configurations [SOTA_4_Backscatter_Multi_Antenna].
@@ VI-B Comparative Analysis with SOTA Baselines
As shown in Fig. 1, BC-TAS significantly outperforms the industry-standard Norm-Based Antenna Selection (NBAS) [SOTA_5_NBAS]. At 10 dB SNR, BC-TAS achieves a BER of $\approx 3.2\times 10^{-4}$, while NBAS remains at $\approx 2.6\times 10^{-2}$. This order-of-magnitude improvement stems from the framework’s ability to exploit frequency-selective diversity across subcarriers rather than relying on a static, wideband norm-based metric. Furthermore, the proposed scheme achieves a near-optimal diversity slope and remains robust even under severe 20% CSI error conditions.
VI-C Impact of CSI Imperfection and Robustness Analysis
Practical backscatter deployments require resilience to CSI estimation errors, which are prevalent due to the low-power nature of the tag. We evaluate the BC-TAS framework under three regimes: Perfect CSI (P-CSI), 5% estimation error, and 20% error. As illustrated in Fig. 1, the P-CSI configuration tracks the MD-OFDM ceiling until 10 dB SNR, effectively exploiting the full spatial-frequency diversity of the TGn Model F environment.
Under moderate imperfection (5% error), BC-TAS exhibits a negligible 1.5 dB SNR penalty at a BER of $10^{-3}$ while maintaining a steep diversity slope. Under severe imperfection (20% error), the system encounters a performance floor at $\approx 1.8\times 10^{-2}$. Despite this, BC-TAS (20% Error) maintains a lower BER than both NBAS and MMSE $\times4$ baselines in the 0–8 dB SNR range. This robustness stems from the selection logic: while NBAS [SOTA_5_NBAS] collapses by erroneously selecting weak paths due to noisy norm estimations, the multi-objective BC-TAS provides a ”smoothing” effect. By considering the aggregate channel state across all 256 subcarriers, BC-TAS minimizes the probability of selecting deep-fade antennas, ensuring link reliability even under high CSI aging or estimation error.
VI-D Computational Efficiency vs. Performance Ceiling
While OES [SOTA_3_Exhaustive_Search] defines the absolute performance ceiling (zero BER at 8 dB), it requires a per-carrier exhaustive search across all $N_{sc}$ subcarriers. As shown in Table II, our BC-TAS heuristic demonstrates a $2.3\times$ speedup over conventional MMSE detectors, completing the simulation in 79.16 seconds compared to 181.69 seconds for MMSE $\times4$. In GPU-accelerated environments, BC-TAS achieves execution times comparable to OES while bridging the gap between theoretical optimality and practical real-time constraints. This latency reduction is vital for energy-constrained WSNs requiring rapid selection to maintain the sensing-communication link.
@@ VI-C Impact of CSI Imperfection and Robustness Analysis
VI-E BER Performance and Environmental Robustness
The diversity gain of BC-TAS is evidenced by the significantly steeper BER slope compared to the MMSE baseline. To validate robustness, performance is analyzed across IEEE 802.11 TGn channel models A–F (Fig. 2).
@@ VI-E BER Performance and Environmental Robustness
While Model A (Flat/Small Office) yields the best performance due to low frequency selectivity, the diversity order remains high even in Model F (Large Open Space). Despite the increased delay spreads and deep fades in larger environments, the MD-OFDM selection diversity achieves a $10^{-3}$ BER at approximately 12–14 dB SNR across all dispersive models. This confirms the framework’s capability to maintain consistent sensing-communication links regardless of the physical indoor layout.
VI-F Outage Probability and Diversity Gain Analysis
To evaluate link reliability in deep-fading indoor environments, we analyze the Cumulative Distribution Function (CDF) of the instantaneous SINR, $\gamma$, in a TGn Model F environment. The outage probability is defined as $P_{\mathrm{out}}=P(\gamma<\gamma_{\mathrm{th}})$, where $\gamma_{\mathrm{th}}=7$ dB represents the minimum threshold for reliable QPSK demodulation.
Figure 3: Outage Probability CDF in TGn Model F at an average $\mathrm{SNR}=16$ dB.
As illustrated in Fig. 3, BC-TAS significantly reshapes the SINR distribution. While conventional random selection tracks the Theoretical Rayleigh (SISO) baseline ($P_{\mathrm{out}}\approx 0.1058$), BC-TAS suppresses the distribution’s lower tail. At the $7$ dB threshold, BC-TAS achieves an outage probability of $1.75\times 10^{-4}$, representing a $604.77\times$ reduction in link outages. As detailed in Table III, this improvement stems from the transition to a diversity order of $L=4$, where reliability gains grow exponentially with $N_{T}$. In Rayleigh fading, $P_{\mathrm{out}}$ scales as $(P_{\mathrm{out,SISO}})^{N_{T}}$; thus, our three-order-of-magnitude reduction aligns with theoretical expectations. By bypassing subcarriers in deep fades, BC-TAS ensures the signal remains above the decoding threshold $\approx 99.99\%$ of the time. Our simulated outage of $1.75\times 10^{-4}$ closely tracks the theoretical floor $(P_{\mathrm{out,SISO}})^{N_{t}}\approx 1.2\times 10^{-4}$, with the slight delta representing the selection penalty $\Delta$ derived in Appendix A.C.
@@ VI-F Outage Probability and Diversity Gain Analysis
VI-G BCF and Link Stability
The stability of the reflected field at the passive tag is governed by the BCF. As derived in Lemma 1, BC-TAS utilizes a field-flattening cost term to minimize temporal peaks. To ensure statistical rigor, we performed $1,000$ Monte Carlo trials per antenna configuration ($N_{t}\in\{2,\dots,32\}$), integrating a scalar Kalman Filter to smooth channel gains and prevent noise-driven antenna switching.
@@ VI-G BCF and Link Stability
As illustrated in Fig. 4, BC-TAS (Kalman) exhibits a superior downward trend in BCF compared to all baselines. At $N_{t}=32$, BC-TAS achieves a BCF of $\approx 14.0$ dB, a $1.1$ dB improvement over OES and $2.5$ dB over NBAS. While OES [SOTA_3_Exhaustive_Search] is optimal for BER, it remains “tag-blind,” often selecting antennas that create constructive interference peaks at the tag. Conversely, BC-TAS intelligently sacrifices marginal primary SNR to maintain a flatter reflected field. The Kalman filter is vital here; by smoothing channel state estimates, it suppresses the high BCF values typically caused by erratic, noise-driven antenna switching across subcarriers. The 1.1 dB BCF advantage over OES confirms Lemma 1, demonstrating that selection diversity can be dual-purposed for link reliability and peak power suppression. Furthermore, the 1,000 Monte Carlo realizations confirm that this is a persistent statistical property of the BC-TAS heuristic rather than a transient benefit.
VI-H PA Distortion Analysis
Hardware impairments are modeled using the Rapp PA ($p=2,V_{sat}=1$), with distortion quantified via EVM and IEEE 802.11n spectral mask compliance to ensure coexistence with legacy narrowband WSN nodes.
@@ VI-H PA Distortion Analysis
As shown in Fig. 5, BC-TAS maintains a consistent linearity advantage over MMSE MU-MIMO. While the baseline remains above the 5.6% EVM limit for 16-QAM throughout most of the transition, BC-TAS achieves compliance at an IBO of approximately 13 dB. This reduction in constellation distortion is complemented by the spectral analysis in Fig. 8. Even at 10 dB IBO, the normalized PSD remains strictly within IEEE 802.11n limits, as selection diversity mitigates the spectral regrowth typical of high-PAPR signals. By operating with less back-off while maintaining signal integrity, BC-TAS maximizes transmit power for tag energy harvesting while ensuring robust coexistence in the 2.4/5 GHz bands.
VI-I Hardware Performance and Complexity Analysis
The hardware feasibility of BC-TAS is evaluated by analyzing its robustness against PA non-linearities and its energy footprint compared to conventional MU-MIMO baselines.
VI-I1 PAPR Reduction and Statistical Analysis
PAPR remains a critical bottleneck for PA efficiency. We evaluate peak power events using the Complementary Cumulative Distribution Function (CCDF), defining the probability that an OFDM symbol’s PAPR exceeds a threshold $\gamma_{0}$.
@@ VI-I1 PAPR Reduction and Statistical Analysis
As shown in Fig. 6, BC-TAS exhibits a profound statistical advantage. At a $10^{-2}$ clipping probability, BC-TAS maintains a PAPR of $\approx$7.1 dB, while the MMSE MU-MIMO baseline reaches 9.5 dB. This 2.4 dB reduction results from the selection-based architecture; by intelligently mapping subcarriers to antennas that minimize constructive interference, BC-TAS effectively “truncates” extreme peak events, driving the improved linearity and energy harvesting observed in subsequent analyses.
VI-I2 Energy Efficiency (EE)
EE is the ratio of spectral efficiency to total power consumption, including RF chain activation and computational costs ($P_{RF}+P_{SEL}$).
@@ VI-I2 Energy Efficiency (EE)
Results in Fig. 7 show BC-TAS provides a massive EE advantage, peaking at 0.075 bits/Hz/mW. Conversely, the MMSE baseline remains below 0.01 bits/Hz/mW due to the simultaneous activation of $N_{t}$ RF chains and high-complexity matrix inversions.
VI-I3 Spectral Mask Compliance
Using the Rapp model ($p=2$), we evaluate spectral regrowth against the IEEE 802.11 mask.
Figure 8: Indoor Spectral Compliance: BC-TAS Transmit Spectrum under IEEE 802.11 Mask.
The normalized PSD in Fig. 8 confirms BC-TAS compliance even at low IBO. Sideband emissions are suppressed below 45-45 dB, allowing the PA to operate nearer saturation to maximize harvestable power for backscatter tags while ensuring coexistence with legacy nodes.
This creates a performance loop: selection diversity reduces PAPR, forcing the signal into the PA’s linear region and liberating power to increase tag harvesting efficiency $\eta_{H}$. By reducing the IBO while maintaining 16-QAM compliance, BC-TAS closes the gap between hardware constraints and passive link power requirements.
VI-J Joint Optimization and Fundamental Trade-off
The BC-TAS framework showcases a fundamental trade-off between primary link reliability (BER) and victim interference suppression ($\Delta I$).